OpenStudy (anonymous):

okay, so the current in a river is 2 miles per hour. A small boat can travel 32 miles upstream and then 24 miles downstream in a total time of 6 hours. find the speed of the boat in still water.

OpenStudy (anonymous):

upstream velocity = velocity of boat(unknown)-velocity of stream down stream = velocity of boat + velocity of stream time = distance/ speed form quadratic eqn and solve

OpenStudy (anonymous):

time = distance/rate Solve the following for r, the rate (speed in still water)\[\frac{32}{r-2}+\frac{32}{r+2}= 6\]

OpenStudy (anonymous):

\[r = \frac{2}{3} \left(8+\sqrt{73}\right) =11.03 \text { mph} \]

OpenStudy (anonymous):

Thank you for the medal. By the way, it was a surprise that, if this was a school problem, x had a radical in it's expression.